Q5.A system of N spin-1/2 particles, each with magnetic moment m, is in presence of a
magnetic field B. Initially, all the magnetic dipoles are oriented along the direction of the
magnetic field. If the energy of the system is increased by ΔE = + mB, what is the change
in entropy of the system?
1
Expert's answer
2020-04-20T19:05:58-0400

As per the given question,

N spin-1/2

magnetic moment of each =m

Magnetic field =B

Increased in the energy of the system=+mB

Change in entropy=?

We know that single bohr frequency "\nu=\dfrac{E_i-E_f}{h}=\dfrac{\omega_o}{2\pi}"

now substituting

"(\nu)=\dfrac{2\pi mB}{h}" N energy

We know that,

"H=\omega_o . S"

For one,

and "H|+>=+\dfrac{h\omega}{2} |+>"

"H|->=-\dfrac{h\omega}{2} |->"

So, change in entropy for N,

"Q= E=mB"


S"=\dfrac{dE}{dT}=\dfrac{NmB}{2}"